# Hunyuan-MT Technical Report

### v1

@@ 3 Post-training Stage
Figure 2: Post-training pipeline of the Hunyuan-MT-7B and Hunyuan-MT-Chemira-7B models.
@@ 4.2 Evaluation Metrics
Hunyuan-MT-Chemira-7B XCOMET-XXL 0.8974 0.8719 0.9306 0.9132 0.8268 0.8787 0.6089
@@ 5.1 Case Study
Table LABEL:table:_test_scenario_1_4, Table LABEL:table:_test_scenario_5_6, and Table LABEL:table:_test_scenario_Chemira illustrate representative translation cases, comparing the outputs of Hunyuan-MT-7B and Hunyuan-MT-Chemira-7B on key testing points. These translation cases demonstrate the model’s sophisticated contextual understanding and proficiency in managing complex translations.
@@
Example #1 Unit 602, HUANGPU Complex, 654 Huangpu Drive, Huangpu District, Guangzhou, Guangdong Province, 510000, China
@@
Building upon this foundation, the Hunyuan-MT-Chemira-7B variant exhibits further refinements in nuanced interpretation. This is exemplified by its ability to resolve ambiguities in specialized domains, such as identifying ”d2” as the game ”Diablo II” and interpreting ”make a game” as a term for in-game trading. The model also shows a sophisticated approach to informal language, pragmatically translating the intensifier ”fucking” to convey emphasis rather than resorting to a literal profanity. In other scenarios, it correctly maps ”三分” to ”three points” in a sports context. It demonstrates greater situational awareness by rendering ”穿过” as ”sped through” a crowd, avoiding the problematic implication of ”drove through.” These cases suggest that the Chemira enhancement facilitates a more fine-grained understanding of colloquialisms and domain-specific context, contributing to translations that are both more accurate and pragmatically sound.
Table 9: Case studies of the Hunyuan-MT-Chemira-7B model.
Scenario 7: Hunyuan-MT-Chemira-7B Case Study
@@
Hunyuan-MT-Chemira-7B I hope we can secure those three points, but it’s going to be a tough match, as they performed well in their last game. However, with so many people supporting us, I’m confident that we can continue to move forward and achieve good results.
@@
Hunyuan-MT-Chemira-7B A 1948 Ford Mercury sped through a crowd of onlookers in the small town of Monroe, Georgia, on its way towards the tiny Moorestown Bridge.
@@ 6 Conclusion
In this report, we introduce Hunyuan-MT-7B and Hunyuan-MT-Chemira-7B, a family of open-source LLMs specifically engineered for machine translation, enabling bidirectional translation across 33 languages. We present a comprehensive overview of our training methodology, which encompasses pre-training, supervised fine-tuning, and reinforcement learning phases. Throughout this exposition, we share critical insights and best practices gleaned from our iterative optimization process. Remarkably, with only 7B parameters, Hunyuan-MT-7B and Hunyuan-MT-Chemira-7B achieve translation quality that rivals—and in certain cases surpasses—that of state-of-the-art LLMs and leading commercial translation systems, as validated through both automatic metrics and human evaluation. By making Hunyuan-MT’s model weights publicly available, we aim to empower the research community with an accessible, high-performance foundation model that can accelerate innovation in MT research and applications.

### v2

@@ 3 Post-training Stage
Figure 2: Post-training pipeline of the Hunyuan-MT-7B and Hunyuan-MT-Chimera-7B models.
@@ 4.2 Evaluation Metrics
Hunyuan-MT-Chimera-7B XCOMET-XXL 0.8974 0.8719 0.9306 0.9132 0.8268 0.8787 0.6089
@@ 5.1 Case Study
Table LABEL:table:_test_scenario_1_4, Table LABEL:table:_test_scenario_5_6, and Table LABEL:table:_test_scenario_Chimera illustrate representative translation cases, comparing the outputs of Hunyuan-MT-7B and Hunyuan-MT-Chimera-7B on key testing points. These translation cases demonstrate the model’s sophisticated contextual understanding and proficiency in managing complex translations.
@@
Example #1 The YMU community mourns the loss of Dr. Wang, a beloved faculty member whose legacy lives on through his students.
Testing Points ”YMU”: Yunnan University of Nationalities.
Huanyuan-MT-7B 云南民族大学(YMU)的师生们为王博士的离世感到悲痛。他是一位深受爱戴的教师,他的精神将通过他的学生们得以延续。
Google-Translator 杨百翰大学全体师生对王博士的去世深感悲痛,王博士是一位受人爱戴的教师,他的精神遗产通过他的学生得以传承。
Example #2 Unit 602, HUANGPU Complex, 654 Huangpu Drive, Huangpu District, Guangzhou, Guangdong Province, 510000, China
@@
Example #3 黄尸癫狗又出黎周围吠人了!
Testing Points Cantonese =¿ Chinese
Huanyuan-MT-7B 黄尸疯狗又出来到处吠人了!
Google-Translator 黄尸癫狗又出黎周围吠人了!
Example #4 [Uncaptioned image]
Testing Points Mongolian =¿ Chinese
Huanyuan-MT-7B 提升“互联网+医疗”水平,内蒙古首家“智慧医院”上线。
Google-Translator [Uncaptioned image]
Example #5 [Uncaptioned image]
Testing Points Uyghur =¿ Chinese
Huanyuan-MT-7B 创新将成为城市繁荣发展的基础。
Google-Translator 创新是城镇化的基础。
Building upon this foundation, the Hunyuan-MT-Chimera-7B variant exhibits further refinements in nuanced interpretation. This is exemplified by its ability to resolve ambiguities in specialized domains, such as identifying ”d2” as the game ”Diablo II” and interpreting ”make a game” as a term for in-game trading. The model also shows a sophisticated approach to informal language, pragmatically translating the intensifier ”fucking” to convey emphasis rather than resorting to a literal profanity. In other scenarios, it correctly maps ”三分” to ”three points” in a sports context. It demonstrates greater situational awareness by rendering ”穿过” as ”sped through” a crowd, avoiding the problematic implication of ”drove through.” These cases suggest that the Chimera enhancement facilitates a more fine-grained understanding of colloquialisms and domain-specific context, contributing to translations that are both more accurate and pragmatically sound.
Table 9: Case studies of the Hunyuan-MT-Chimera-7B model.
Scenario 7: Hunyuan-MT-Chimera-7B Case Study
@@
Hunyuan-MT-
Chimera-7B
I hope we can secure those three points, but it’s going to be a tough match, as they performed well in their last game. However, with so many people supporting us, I’m confident that we can continue to move forward and achieve good results.
@@
Hunyuan-MT-
Chimera-7B
A 1948 Ford Mercury sped through a crowd of onlookers in the small town of Monroe, Georgia, on its way towards the tiny Moorestown Bridge.
@@ 6 Conclusion
In this report, we introduce Hunyuan-MT-7B and Hunyuan-MT-Chimera-7B, a family of open-source LLMs specifically engineered for machine translation, enabling bidirectional translation across 33 languages. We present a comprehensive overview of our training methodology, which encompasses pre-training, supervised fine-tuning, and reinforcement learning phases. Throughout this exposition, we share critical insights and best practices gleaned from our iterative optimization process. Remarkably, with only 7B parameters, Hunyuan-MT-7B and Hunyuan-MT-Chimera-7B achieve translation quality that rivals—and in certain cases surpasses—that of state-of-the-art LLMs and leading commercial translation systems, as validated through both automatic metrics and human evaluation. By making Hunyuan-MT’s model weights publicly available, we aim to empower the research community with an accessible, high-performance foundation model that can accelerate innovation in MT research and applications.